OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1). parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (anonymous):

That looks like the 3rd choice because you want 1 for the slope of the parallel line and -1 for the slope of the perpendicular line. And I checked the points.

OpenStudy (anonymous):

Both equations are first gotten through the point-slope formula for the equation of a line:\[y - y _{1} = m(x - x _{1})\]where "m" is your slope.

OpenStudy (anonymous):

So, for the parallel line: y - (-1) = (1)(x - (-2)) -> y = x + 1

OpenStudy (anonymous):

what answer is that

OpenStudy (anonymous):

For the perpendicular line: y - (-1) = (-1)(x - (-2)) -> y = -x - 3 And the overall answer is the 3rd one.